OpenStudy (anonymous):

A Honda Civic travels in a straight line along a road. Its distance x from a stop sign is given as a function of time t by the equation x(t)= α t2− β t3, where α = 1.42m/s2 and β = 5.00×10−2m/s3 . Calculate the average velocity of the car for the time interval t=0 to t1 = 2.08s . Need help Asap

OpenStudy (anonymous):

dude..... α is in m/s2 as acceleration..... i got it...... but what happen to the β which is in m/s3.... which unit is it...... clarity please..........

OpenStudy (anonymous):

idk im really confused at thsi question

OpenStudy (anonymous):

@harishk see whole term β t3 which is unit of x

OpenStudy (anonymous):

x(t)= α t2− β t3 calculate v(t)= dx(t)/dt =x(t)= 2α t− 3β t2 so now calculate v(t) at t=0 and t1 = 2.08s now average velocity would be v(average)= [v(t1 = 2.08s)-v(t=0)] / (2.08-0)

OpenStudy (anonymous):

sorry in v(t)= dx(t)/dt =x(t) i wrote x(t) by mistake

OpenStudy (anonymous):

I understand the the first 2 steps but what do u have to do in the end sorry

OpenStudy (anonymous):

that is gyanu its good

OpenStudy (anonymous):

oh @thereyruls sorry i made huge mistake average velocity=( distance travel by car in time interval t=0 to t1 = 2.08)/time interval = {x(2.08)-x(0)}/(2.08-0) where x(2.08) is value of x(t) at t=2.08 in x(t)= α t2− β t3, put t= 2.08 you will get x(2.08) similarly in x(t)= α t2− β t3, put t=0 you will get x(0) please tell me if you don't get any thing and lots of sorry for my mistake.